The base of a ladder is 15 feet away from the wall. The top of the ladder is 16 feet from the floor. Find the length of the ladder to the nearest thousandth.
20.911 feet
21.932 feet
19.800 feet
24.233 feet

Answer by nerdybill(7384)

The base of a ladder is 15 feet away from the wall. The top of the ladder is 16 feet from the floor. Find the length of the ladder to the nearest thousandth.
20.911 feet
21.932 feet
19.800 feet
24.233 feet
.
Apply Pythagorean theorem:
Let x = length of the ladder
then
x^2 = 15^2 + 16^2
x^2 = 225 + 256
x^2 = 481
x = 21.932 feet